OpenStudy (anonymous):

Can someone show me how one would find if The sum of 1/3^n would converge? I don't know what test to use...

OpenStudy (anonymous):

I'm thinking comparison test, but I don't know how to show anything besides p-test

OpenStudy (anonymous):

\[\sum_{n=1}^{\inf} 1/3^n\]

Directrix (directrix):

Here is what I am thinking:

OpenStudy (anonymous):

ohhhh a geometric sequence! so you are saying the geometric series is 3^(-n)? I can use that to compare it to.

Directrix (directrix):

The first term would be 1/3, right? So, a = 1/3 The common ratio, r = also 1/3.

OpenStudy (anonymous):

=1/2, satisfying |r|<1, Thanks!

Directrix (directrix):

3^(-n) is the same as (1/3)^n So, I did 1/3 /(1 - (1/3) ) = 1/2 which is what you got.